elimination method for solving A+8B=12 A+3B=2
Answer by ankor@dixie-net.com(22740)

elimination method for solving
A + 8B = 12
A + 3B = 2
----------------subtraction eliminates A find B
0 + 5B = 10
B = 10/5
B = 2
:
Use the 1st equation to find A, replace B with 2
A + 8(2) = 12
A + 16 = 12
A = 12 - 16
A = -4
:
:
Check solutions in the 2nd equation
A + 3B = 2
Replace A and B
-4 + 3(2) = 2
-4 + 6 = 2